# and Efficiency

### v1

@@ Abstract
SQL query rewriting aims to reformulate a query into a more efficient form while preserving equivalence. Most existing methods rely on predefined rewrite rules. However, such rule-based approaches face fundamental limitations: (1) fixed rule sets generalize poorly to novel query patterns and struggle with complex queries; (2) a wide range of effective rewriting strategies cannot be fully captured by declarative rules. To overcome these issues, we propose using large language models (LLMs) to generate rewrites. LLMs can capture complex strategies, such as evaluation reordering and CTE rewriting. Despite this potential, directly applying LLMs often results in suboptimal or non-equivalent rewrites due to a lack of execution awareness and semantic grounding. To address these challenges, We present E3-Rewrite, an LLM-based SQL rewriting framework that produces executable, equivalent, and efficient queries. It integrates two core components: a context construction module and a reinforcement learning framework. First, the context module leverages execution plans and retrieved demonstrations to build bottleneck-aware prompts that guide inference-time rewriting. Second, we design a reward function targeting executability, equivalence, and efficiency, evaluated via syntax checks, equivalence verification, and cost estimation. Third, to ensure stable multi-objective learning, we adopt a staged curriculum that first emphasizes executability and equivalence, then gradually incorporates efficiency. Extensive experiments show that E3-Rewrite achieves up to a 25.6% reduction in query execution time compared to state-of-the-art methods across multiple SQL benchmarks. Moreover, it delivers up to 24.4% more successful rewrites, expanding coverage to complex queries that previous systems failed to handle.
@@ Introduction
More fundamentally, rule-based rewriting systems suffer from limitations that restrict their expressiveness, composability, and adaptability. First, fixed rule sets lack expressiveness and struggle to capture structural rewrites beyond predefined patterns (Begoli et al. 2018). Second, rule dependencies form an entangled and fragile search space, posing challenges for both traditional search-based methods (Zhou et al. 2021) and LLM-driven prompting (Li et al. 2024), often resulting in brittle recommendations. Third, effective rewriting strategies inherently fall outside the expressive capacity of declarative rules (Dong et al. 2023; Leis et al. 2015), especially those involving Common Table Expressions (CTEs) or requiring fine-grained control over evaluation order and sub-plan reuse (Wang et al. 2022). Most importantly, existing methods cannot adapt to new query structures or workloads over time, as they lack mechanisms to learn from execution feedback and refine their rewriting strategies.
@@
Query Rewriting. Existing query rewriting methods have evolved from rule-based systems to learning-based and LLM-augmented approaches.(1) Heuristic Rule-based Systems. Early systems like PostgreSQL (PostgreSQL 2025) and Volcano (Graefe and McKenna 1993) rely on fixed or heuristic rule sequences. These approaches often neglect inter-rule dependencies, leading to blind or inefficient search. (2) Learning-based Rule Selection. LearnedRewrite (Zhou et al. 2021) treats rewriting as a rule search problem and uses MCTS with learned cost models. However, generalization across schemas remains limited due to reliance on schema-specific reward estimators. (3) LLM-Augmented Rewriting. LLM-R² (Li et al. 2024) and R-Bot (Sun, Zhou, and Li 2024) leverage LLMs for rule retrieval or recommendation via prompting. While promising, they still face issues in semantic alignment, cost feedback, and execution awareness. These limitations motivate the need for a fully learnable, execution-driven rewriting framework that generalizes beyond rule enumeration.

### v2

@@ Abstract
SQL query rewriting aims to reformulate a query into a more efficient form while preserving equivalence. Most existing methods rely on predefined rewrite rules. However, such rule-based approaches face fundamental limitations: (1) fixed rule sets generalize poorly to novel query patterns and struggle with complex queries; (2) a wide range of effective rewriting strategies cannot be fully captured by declarative rules. To overcome these issues, we propose using large language models (LLMs) to generate rewrites. LLMs can capture complex strategies, such as evaluation reordering and CTE rewriting. Despite this potential, directly applying LLMs often results in performance regressions or non-equivalent rewrites due to a lack of execution awareness and semantic grounding. To address these challenges, We present E3-Rewrite, an LLM-based SQL rewriting framework that produces executable, equivalent, and efficient queries. It integrates two core components: a context construction module and a reinforcement learning framework. First, the context module leverages execution plans and retrieved demonstrations to build bottleneck-aware prompts that guide inference-time rewriting. Second, we design a reward function targeting executability, equivalence, and efficiency, evaluated via syntax checks, equivalence verification, and cost estimation. Third, to ensure stable multi-objective learning, we adopt a staged curriculum that first emphasizes executability and equivalence, then gradually incorporates efficiency. Across multiple SQL benchmarks, our experiments demonstrate that E3-Rewrite can shorten query execution time by as much as 25.6% relative to leading baselines, while also producing up to 24.4% more rewrites that meet strict equivalence criteria. These gains extend to challenging query patterns that prior approaches could not effectively optimize.
@@ Introduction
More fundamentally, rule-based rewriting systems suffer from limitations that restrict their expressiveness, composability, and adaptability. First, fixed rule sets lack expressiveness and struggle to capture structural rewrites beyond predefined patterns (Begoli et al. 2018; Liu and Mozafari 2024). Second, rule dependencies form an entangled and fragile search space, posing challenges for both traditional search-based methods (Zhou et al. 2021) and LLM-driven prompting (Li et al. 2024), often resulting in brittle recommendations. Third, effective rewriting strategies inherently fall outside the expressive capacity of declarative rules (Dong et al. 2023; Leis et al. 2015), especially those involving Common Table Expressions (CTEs) or requiring fine-grained control over evaluation order and sub-plan reuse (Wang et al. 2022). Most importantly, existing methods cannot adapt to new query structures or workloads over time, as they lack mechanisms to learn from execution feedback and refine their rewriting strategies. A recent concurrent approach, QUITE (Song et al. 2025), explores a training-free, feedback-aware multi-agent framework that iteratively refines rewrites at inference time via a hybrid SQL corrector and hint injection. While effective in certain settings, such inference-time correction strategies may require multiple iterations and lack integrated training-time feedback, making it challenging to jointly optimize executability, equivalence, and efficiency.
@@
Query Rewriting. Existing query rewriting methods have evolved from rule-based systems to learning-based and LLM-augmented approaches.(1) Heuristic Rule-based Systems. Early systems like PostgreSQL (PostgreSQL 2025) and Volcano (Graefe and McKenna 1993) rely on fixed or heuristic rule sequences. These approaches often neglect inter-rule dependencies, leading to blind or inefficient search. (2) Learning-based Rule Selection. LearnedRewrite (Zhou et al. 2021) treats rewriting as a rule search problem and uses MCTS with learned cost models. However, generalization across schemas remains limited due to reliance on schema-specific reward estimators. (3) LLM-Augmented Rewriting. LLM-R² (Li et al. 2024) and R-Bot (Sun, Zhou, and Li 2024) leverage LLMs for rule retrieval or recommendation via prompting. QUITE (Song et al. 2025) adopts a training-free, feedback-aware multi-agent pipeline that combines a structured knowledge base, hybrid SQL corrector, and hint injection to iteratively refine rewrites. While promising, such inference-time correction frameworks often struggle to simultaneously ensure semantic equivalence and achieve performance gains, and may require multiple iterative inference rounds to converge on a satisfactory rewrite, which significantly increases execution time. These limitations motivate the need for a fully learnable, execution-driven rewriting framework that generalizes beyond rule enumeration.